Tuesday 03/09/16

9:45 am – 7:15 pm

Probe synthesis

Set up RT reactions

  • ran RT reactions
  • ran gel
  • ran oligo cleanup
  • see updated protocol
  • Gel results: (Lib7-10 kb BX-C, Lib7-2kb BX-C, Lib8-15 kb en), DNA left, Cy3 right (with gel green blead-through).

Project planning

  • prepare slides for meeting with XZ discuss next chromatin project
  • discuss project design with BB
  • meeting with XZ

Orders & Paperwork

  • place T7 order with NEB
  • sent request to Alec for PolyD Lysine order
  • send paperwork to UCSF
This entry was posted in Summaries. Bookmark the permalink.